# Know Your Neighborhood: General and Zero-Shot Capable Binary Function Search Powered by Call Graphlets

### v1

@@ 4.5 Mono-Architecture Binary Function Search (x86-64) (RQ2)
The results from these experiments are in Table 7. Both of our model variants out-perform the baseline approaches across both of the reported search pool sizes of 100 and 10000. Our KYN-Cross model performs the strongest within the 100 search pool size with a relative percentage increase of 23.4% R@1 and 12.2% MRR@10 when compared against jTrans and a relative percentage increase of 6.1% R@1 and 2.6% MRR@10 when compared against HermesSim. The difference in performance increases significantly when looking at the large search pool size of 10,000. Both model variants significantly outperform both of the benchmark approaches with the relative percentage increases of 53.95% (KYN-Cross) / 55.29% (KYN-x8664) R@1 and 46.25% (KYN-Cross) / 44.97% (KYN-x8664) MRR@10 when compared against jTrans. When compared against HermesSim, the performance increase is less but still significant with a relative percentage increase of 12.66% (KYN-Cross) / 14.11% (KYN-x8664) R@1 and 9.26% (KYN-Cross) / 7.92% (KYN-x8864).
@@
100 10000
jTrans 65.0/73.8 31.4/37.4
HermesSim 75.6/80.7 48.1/54.6
KYN-Cross 80.2/82.8 54.6/59.9
KYN-x8664 77.2/80.4 55.4/59.1
Table 7: x86-64 Mono-Architecture results on Cisco Dataset-1. Results are presented with R@1 first and then MRR@10.
To further evaluate our model and understand its performance when faced with a larger x86-64 dataset, we also evaluated it against BinaryCorp-3M with and without duplicates. These results can be seen in Table 8 and show that both model variants perform similarly with the cross-architecture variant typically edging ahead. There are several areas of note. Firstly, the performance of the models when faced with a different mono-architecture dataset is significantly degraded. This highlights the need to evaluate against a range of different dataset to get a true understanding of a models performance. Secondly, the performance is inflated significantly when duplicates are present across both search pool sizes tested. This suggests that the inclusion of duplicates within evaluation datasets is a critical mistake and may lead to misrepresentation of the quality and performance of approaches.
@@
100 10000 100 10000
KYN-Cross 68.0/73.9 38.8/44.1 51.6/58.8 31.2/36.9
KYN-x86-64 67.0/73.6 36.6/41.9 51.8/58.0 31.4/36.2
Table 8: Mono-Architecture Binary Function Search Results. WD denotes "with duplicates" whilst ND denotes "no duplicates"
@@
Table 9: Zero Shot Vulnerability Search Results
@@ 4.6 Cross-Architecture Binary Function Search with Inlining (RQ3)
The following section evaluates KYN in the presence of the potential of function inlining. Function inlining introductions the possibility for callee functions to be merged into (or inlined) into a caller function. If this occurs, the structure of the caller function can change significantly. This is an area which is typically highlighted as a limitation of approaches or something that is explicitly turned off [19, 15]. We chose instead of tackle this area head-on and evaluate our approach against a BinKit variant called BinKit-Normal. Within this dataset, all of the binaries have been compiled where there is a possibility for function inlining. The results presented below in Table 10 show the metric value for a given search pool size as well as the percentage difference when compared against the results present for BinkitNoInline in Table 6.
@@
100 0.73 11.6% 0.79 9.3% 0.82 8.17%
250 0.66 14.99% 0.72 13.36% 0.75 12.47%
1000 0.59 7.39% 0.66 7.42% 0.69 7.86%
10000 0.33 23.39% 0.42 20.23% 0.47 18.47%
Table 10: Binkit-Normal Metrics Across Different Search Pools with comparison against Binkit-NoInline. The percentages represent percentage decrease when function inlining was present.
The key observations from the results are as follows. Across all search pool sizes, there is a performance degradation of between 7% to 23.5% or in other works between 0.07 and 0.1 in absolute terms across the reported metrics. These results do show that function inlining does have a negative impact on our methods performance but is not dramatic. Considering that our approach was trained with a dataset that does not include function inlining , the results show our model has learnt general representation allowing it to maintain reasonable performance when faced with function inlining.
@@ 4.7 Zero-Shot Cross-Architecture Vulnerability Search (RQ4)
We also conduct a zero-shot experiment using the same methodology as the N-day vulnerability search task described in [15] but compile the same libcrypto.so library for RISC-V and PowerPC 32-bit to evaluate our approaches model within a zero-shot context with the results reported in Table 9.
The performance within a zero-shot context area mixed depending on the architecture used for search pool functions (i.e the architecture of what is being searched). Focusing on the ARM32 firmware image, KYN demonstrates high accuracy and precision in identifying OpenSSL vulnerabilities when provided with ARM32, x86-64, and x86-32 query functions, evidenced by a mean rank of 1 and median rank of 1. However, when faced with a MIPS32 query function, the model’s performance degrades, with a mean rank of 91 and a median rank of 8. Regarding zero-shot architectures, the RISCV-32 performance is acceptable, with a mean rank of 82 and a median rank of 1, albeit skewed by a single poor retrieval. Conversely, the PowerPC results are significantly worse, with a mean rank of 309, a median rank of 134, and a particularly high ranking of 967, indicating high variance in the model’s performance depending on the function encountered.
In addition to the ARM32 firmware image, we also report the results for the MIPS32 image within the same dataset. This is something that other works have typically omitted. Examining the MIPS32 architecture firmware, the results with a MIPS32 query function are satisfactory but inferior to the ARM32 vs. ARM32 results, with a mean score of 53 and a median of 2. When considering the in-domain architectures of ARM32, X86-64, and X86-32, the model performs acceptably, with mean ranks ranging from 57-69 and median ranks from 44-75.5, albeit not as well as the ARM32 vs. In-Domain experiments. Notably, for the out-of-domain RISCV-32 architecture, the mean and median ranks outperform all in-domain architectures. However, consistent with the ARM32 results, PowerPC32 performs poorly, with a mean rank of 433 and a median rank of 479.
The results are mixed but show promise when compared against those reported within [4], especially for zero-shot performance of RISCV-32 $\to$ ARM32. Our findings are however consistent with [4] in terms of MIPS32 being inherently more difficult and challenging that other architectures. The reasons for this are unclear and may warrant additional research.
@@ 4.8 Ablation Study (RQ5)
In order to understand the performance of our model in more detail and the impact of our design decisions, we trained several models on the same dataset and then evaluated them on the same test set, utilizing a search pools of 100 for evaluation generate empirical evidence. KYN is the model architecture used within the rest of the paper above, KYN-NE is the same model architecture but the input graphs have no edge weights and therefore the edge support for GraphConv layers was not used and KYN-NES further amended whereby we replace the softmax aggregation function for add within each of the convolutional layers. The results for each of these settings can be seen below in Table 11.
@@
Table 11: Metric scores across search pools of size 100
The results demonstrate that both the softmax aggregation and edge features contribute to increasing performance. The softmax aggregation provides a modest performance absolute improvement of 0.01 for both MRR@10 and R@1. The edge feature information however contributes to a significant absolute improvement of 0.04 MRR@10 and 0.05 of R@1. These results suggest that incorporating edge weights, in our case edge betweenness, increases the quality of the function representations for graph neural network based binary code similarity detection tasks.
During the experimentation and model development for this study, several experiments were run to understand the usefulness of various types of graph normalization techniques. Namely Layer Normalization [1], GraphNorm Normalization [3] and Instance Normalization [22]. To understand the effect of each, we trained the same model with the same train and test set but changed the normalization technique used within each of the 3 GNN layers.
@@
Table 12: Metric cross across Normalization Choices
The results presented for the normalization experiments in Table 12 are clear. They suggest that, for our data at least, the chosen normalization approach has a profound impact of the overall performance. When comparing GraphNorm to Instance Normalization, the impact of choosing GraphNorm over Instance Normalization is approximately a 0.3 absolute difference or a 42% relative difference. The effects are marginally less for LayerNorm but still significant with approximately a 0.17 absolute difference or 25% relative difference when compared against InstanceNorm.
The reason for this large performance difference likely lies within the large variations, in particular with the number of instructions within our chosen coarse node feature vectors and how both LayerNorm and GraphNorm apply normalization. Both of these normalization approaches seek to learn a normalization function for batches of graphs. Given that each of our batches contain graphs sampled from across both architectures and software projects, the variation within the batches is likely to be high. During training, the normalization functions are trained and this variation is normalized as expected. When moving to test however, when the model is faced with new software projects whereby the call graphlets have a completely different set of nodes and therefore distribution, when the learnt normalization function is applied, the normalization performs poorly. This is especially true when you look at the binaries contained within the train and test sets of the Dataset-1 where the model is initially faced with several modestly sized projects during training before then being faced with a test dataset dominated by a very large and complex project (z3). This issue however does not effect InstanceNorm because the normalization is applied to each graph independently within a batch.
@@ 5 Limitations
Another limitation of the current approach is its reliance on accurate disassembly and call graph recovery. While the methodology employed in this work assumes reliable disassembly and call graph extraction, real-world scenarios may involve obfuscated or heavily optimized binaries, which could impede accurate recovery of these structures. Techniques such as control flow flattening, virtualization obfuscation, or extreme inlining could potentially undermine the effectiveness of the proposed approach. To address this limitation, future research could investigate incorporating deobfuscation strategies or developing more robust graph recovery mechanisms that can handle adversarial obfuscation techniques.
Furthermore, the current evaluation framework primarily focuses on open-source software and benchmark datasets. While these datasets provide a valuable foundation for assessing the approach’s performance, they may not fully capture the diversity and complexity present in real-world binary analysis scenarios, such as those encountered in embedded software reverse engineering or malware analysis settings. Future work should aim to expand the evaluation scope by incorporating a broader range of proprietary software, firmware, and malware samples to further validate the approach’s generalizability and robustness under diverse and potentially adversarial conditions.
@@ Acknowledgments
I would like to acknowledge the significant contribution of the authors of [15] who have provided a foundation for robust comparison within this domain and stimulated significant research interest.
Availability
All components of this paper are publicly available. The training dataset, model training and evaluation code as well as the model artifacts themselves can be found http://repository-to-be-made. The tool used to generate the dataset is also open source and can be found here https://github.com/br0kej/bin2ml.

### v2

@@ 4.5 Mono-Architecture Binary Function Search (x86-64) (RQ2)
The results from these experiments are in Table VII. Both of our model variants out-perform the baseline approaches across both of the reported search pool sizes of 100 and 10,000. Our KYN-Cross model performs the strongest within the 100 search pool size with a relative percentage increase of 19% R@1 and 11% MRR@10 when compared against jTrans and a relative percentage increase of 6% R@1 and 3% MRR@10 when compared against HermesSim. The difference in performance increases significantly when looking at the large search pool size of 10,000. Both model variants significantly outperform both of the benchmark approaches with the relative percentage increases of 52% (KYN-Cross) / 53% (KYN-x8664) R@1 and 38% (KYN-Cross) / 37% (KYN-x8664) MRR@10 when compared against jTrans. When compared against HermesSim, the performance increase is less but still significant with a relative percentage increase of 12% (KYN-Cross) / 13% (KYN-x8664) R@1 and 9% (KYN-Cross) / 7% (KYN-x8864).
@@
100 10,000
jTrans 0.65/0.738 0.314/0.374
HermesSim 0.756/0.807 0.481/0.546
KYN-Cross 0.802/0.828 0.546/0.599
KYN-x8664 0.772/0.804 0.554/0.591
TABLE VII: x86-64 Mono-Architecture results on Cisco Dataset-1. Results are presented with R@1 first and then MRR@10.
To further evaluate our model and understand its performance when faced with a larger x86-64 dataset, we also evaluated it against BinaryCorp-3M with and without duplicates. These results can be seen in Table VIII and show that both model variants perform similarly with the cross-architecture variant typically edging ahead. There are several areas of note. Firstly, the performance of the models when faced with a different mono-architecture dataset is significantly degraded. This highlights the need to evaluate against a range of different dataset to get a true understanding of a models performance. Secondly, the performance is inflated when duplicates are present across both search pool sizes tested. This suggests that the inclusion of duplicates within evaluation datasets is a critical mistake and may lead to misrepresentation of the quality and performance of approaches.
@@
100 10,000 100 10,000
KYN-Cross 0.672/0.735 0.388/0.441 0.624/0.694 0.312/0.369
KYN-x86-64 0.67/0.756 0.366/0.419 0.518/0.58 0.314/0.362
TABLE VIII: Mono-Architecture Binary Function Search Results. R@1 and MMR@10. WD denotes ”with duplicates” whilst ND denotes ”no duplicates”
@@
4.5.1 Summary
The results above answer RQ2 and demonstrate that our KYN architecture model outperforms the baseline approaches significantly when faced with unseen libraries, optimization options and compilers within a mono-architecture setting. We also conduct an additional experiment that highlights the impact duplicate functions can have on evaluation performance and hope that these findings will be used to enhance future research methodologies.
@@ 4.5.1 Summary
TABLE IX: Zero Shot Vulnerability Search Results
@@ 4.6 Cross-Architecture Binary Function Search with Inlining (RQ3)
The following section evaluates KYN in the presence of the potential of function inlining. Function inlining introductions the possibility for callee functions to be merged into (or inlined) into a caller function. If this occurs, the structure of the caller function can change significantly. This is an area which is typically highlighted as a limitation of approaches or something that is explicitly turned off [11, 12]. We chose instead of tackle this area head-on and evaluate our approach against a BinKit variant called BinKit-Normal. Within this dataset, all the binaries have been compiled where there is a possibility for function inlining. The results presented below in Table X show the metric value for a given search pool size as well as the percentage difference when compared against the results present for BinkitNoInline in Table VI.
@@
100 0.722 14.7% 0.781 11.9% 0.809 10.6%
250 0.648 16.6% 0.716 14.1% 0.749 12.9%
1,000 0.569 14.8% 0.620 14.8% 0.648 14.6%
10,000 0.317 28.3% 0.402 24.7% 0.456 21.7%
TABLE X: Binkit-Normal Metrics Across Different Search Pools with comparison against Binkit-NoInline. The percentages represent percentage decrease when function inlining was present.
The key observations from the results are as follows. Across all search pool sizes, there is a performance degradation of between 10% to 28.3% with an average performance degradation of 16.6%. These results do show that function inlining does have a negative impact on our method’s performance but is not dramatic. Considering that our approach was trained with a dataset that does not include function inlining at all, the results show our model has learned general representation allowing it to maintain reasonable performance when faced with function inlining. This is something that is typically not explored within other works.
4.6.1 Summary
The results above answer RQ3 and show that our model is fairly robust to function inlining even though it was trained on a dataset which does not include inlined examples. We attribute this to our novel call graphlet input data structure as well as the models’ architecture which enable the model to learn general representations of function neighborhoods.
@@ 4.7 Zero-Shot Cross-Architecture Vulnerability Search (RQ4)
We also conduct a zero-shot experiment using the same methodology as the N-day vulnerability search task described in [12] but compile the same libcrypto.so library for RISC-V and PowerPC 32-bit to evaluate our approaches model within a zero-shot context with the results reported in Table IX.
The performance within a zero-shot context are mixed depending on the architecture used for search pool functions (i.e the architecture of what is being searched). Focusing on the ARM32 firmware image, KYN demonstrates good performance in identifying OpenSSL vulnerabilities when provided with ARM32, x86-64, and x86-32 query functions, evidenced by a mean rank of 1 and median rank of 1. However, when faced with a MIPS32 query function, the model’s performance degrades, with a mean rank of 91 and a median rank of 8. Regarding zero-shot architectures, the RISCV-32 performance is acceptable, with a mean rank of 82 and a median rank of 1, albeit skewed by a single poor retrieval. Conversely, the PowerPC results are significantly worse, with a mean rank of 309, a median rank of 134, and a particularly high ranking of 967, indicating high variance in the model’s performance depending on the architecture encountered.
In addition to the ARM32 firmware image, we also report the results for the MIPS32 image within the same dataset. This is something that other works have typically omitted. Examining the MIPS32 architecture firmware, the results with a MIPS32 query function are satisfactory but inferior to the ARM32 vs ARM32 results, with a mean score of 53 and a median of 2. When considering the in-domain architectures of ARM32, X86-64, and X86-32, the model performs acceptably, with mean ranks ranging from 57-69 and median ranks from 44-75.5, albeit not as well as the ARM32 vs In-Domain experiments. Notably, for the out-of-domain RISCV-32 architecture, the mean and median ranks outperform all in-domain architectures. However, consistent with the ARM32 results, PowerPC32 performs poorly, with a mean rank of 433 and a median rank of 479.
4.8 Summary
The results above provide mixed evidence for RQ4. The results show promise when compared against those reported within [5], especially for zero-shot performance of RISCV-32 $\to$ ARM32 but do not demonstrate general zero-shot performance across all of those architectures tested. Our findings are however consistent with Collyer [5] in terms of MIPS32 being inherently more difficult and challenging that other architectures. The reasons for this are unclear and may warrant additional research.
@@ 4.9 Ablation Study (RQ5)
In order to understand the performance of our model in more detail and the impact of our design decisions, we trained several models on the same dataset and then evaluated them on the same test set, utilizing a search pools of 100 for evaluation generate empirical evidence. KYN is the model architecture used within the rest of the paper above, KYN-NE is the same model architecture, but the input graphs have no edge weights and therefore the edge support for GraphConv layers was not used and KYN-NES further amended whereby we replace the softmax aggregation function for add within each of the convolutional layers. The results for each of these settings can be seen below in Table XI.
@@
TABLE XI: Metric scores across search pools of size 100
The results demonstrate that both the softmax aggregation and edge features contribute to increasing performance. The softmax aggregation provides a modest performance absolute improvement of 0.01 for both MRR@10 and R@1. The edge feature information however contributes to a significant absolute improvement of 0.04 MRR@10 and 0.05 of R@1. These results suggest that incorporating edge weights, in our case edge-betweenness, increases the quality of the function representations for graph neural network based binary code similarity detection tasks.
During the experimentation and model development for this study, several experiments were run to understand the usefulness of various types of graph normalization techniques. Namely, Layer Normalization [26], GraphNorm Normalization [27] and Instance Normalization [20]. To understand the effect of each, we trained the same model with the same train and test sets but changed the normalization technique used within each of the 3 GNN layers.
@@
TABLE XII: Metric cross across Normalization Choices
The results presented for the normalization experiments in Table XII are clear. They suggest that, for our training data at least, the chosen normalization approach has a profound impact of the overall performance. When comparing GraphNorm to Instance Normalization, the impact of choosing GraphNorm over Instance Normalization is approximately a 0.3 absolute difference or a 42% relative difference. The effects are marginally less for LayerNorm but still significant with approximately a 0.17 absolute difference or 25% relative difference when compared against InstanceNorm.
The reason for this large performance difference likely lies within the large variations present within our data, in particular with the number of instructions within our chosen coarse node feature vectors and how both LayerNorm and GraphNorm apply normalization. Both of these normalization approaches seek to learn a normalization function for batches of graphs. Given that each of our batches contain graphs sampled from across both architectures and software projects, the variation within the batches is likely to be high. During training, the normalization functions learnable parameters are optimized to reduce this variance leading to improved model performance. When moving to the test set, the feature distribution changes due to being faced with software binaries which are out of distribution. This means that when the normalization is applied, it leads to poor performance. This is especially true when you look at the binaries contained within the train and test sets of the Dataset-1 where the model is initially faced with several modestly sized projects during training before then being faced with a test dataset dominated by a very large and complex project (z3). This issue however does not affect Instance Norm because the normalization is applied to each graph independently within a batch.
4.10 Summary
In terms of answering RQ5, the combination of Instance Normalisation, Softmax Aggregation and edge weights all contribute to the performance of our approach. Our first ablation study clearly show that the inclusion of edge weights provides the biggest performance increase with a modest increase provided by the softmax aggregation. The findings from our second ablation study show a dramatic impact of normalization choice which we attribute to the feature variance across different architectures and software libraries.
@@ 5 Limitations
Another limitation of the current approach is its reliance on accurate disassembly and call graph recovery. While the methodology employed in this work assumes reliable disassembly and call graph extraction, real-world scenarios may involve obfuscated or heavily optimized binaries, which could impede accurate recovery of these structures. Techniques such as control flow flattening, virtualization obfuscation and extreme inlining could potentially undermine the effectiveness of the proposed approach. To address this limitation, future research will investigate incorporating deobfuscation strategies and developing more robust graph recovery mechanisms that can handle adversarial obfuscation techniques.
The current evaluation framework primarily focuses on open-source software and benchmark datasets. While these provide a valuable foundation for assessing the approach’s performance, they may not fully capture the diversity and complexity present in real-world binary analysis scenarios, such as those encountered in embedded software reverse engineering or malware analysis settings. Future work should aim to expand the evaluation scope by incorporating a broader range of proprietary software, firmware, and malware samples to further validate the approach’s generalizability and robustness under diverse and potentially adversarial conditions.
@@ Acknowledgments
I would like to acknowledge the significant contribution of the Marcelli et al [12] who have provided a foundation for robust comparison within this domain and stimulated significant research interest.